ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 250, 2022

Aguiar, M. P. S. and P. A. Zalesskii The profinite completion of the fundamental group of infinite graphs of groups

429

Aomoto, K. See Y. Fedorov Baranov, A., Y. Belov and A. Kulikov Spectral synthesis for

exponentials and logarithmic length

403

Bayart, F., F. Costa Jr. and D. Papathanasiou Disjoint and common hypercyclic algebras

211

Belov, Y. See A. Baranov Cossey, J. and Y. Yang A bound on the diameter of the Brauer

graph of a block of a solvable group

115

Costa Jr., F. See F. Bayart Eckhardt, J., A. Kostenko and T. Kukuljan On the absolutely

continuous spectrum of generalized indefinite strings II

307

Fedorov, Y., J. Komeda, S. Matsutani, E. Previato and K. Aomoto The sigma function over a family of curves with a singular fiber

345

K _{}akol, J., W. Marciszewski, D. Sobota and L. Zdomskyy On complemented copies of the space c_{0} in spaces C_{p}(X × Y)

139

Kaliman, S. Lines in affine toric varieties  85

Komeda, J. See Y. Fedorov Kostenko, A. See J. Eckhardt Kukuljan, T. See J. Eckhardt Kulikov, A. See A. Baranov Marciszewski, W. See J. K _{}akol Matsutani, S. See Y. Fedorov McInerney, M. and K. M. Ng Multiple genericity: a new transfinite

hierarchy of genericity notions

1

Ng, K. M. See M. McInerney Papathanasiou, D. See F. Bayart Park, J.-H. and M. Powell A ribbon obstruction and derivatives

of knots

265

Powell, M. See J.-H. Park Previato, E. See Y. Fedorov Saienko, M. Characterisation of valuations and curvature measures

in euclidean spaces

463

Sarkar, J. On the pointwise converse of Fatou’s theorem for Euclidean and real hyperbolic spaces

179

Sobota, D. See J. K _{}akol Thomas, J. Delocalisation of eigenfunctions on large genus random

surfaces

53

Yang, Y. See J. Cossey Zalesskii, P. A. See M. P. S. Aguiar Zdomskyy, L. See J. K _{}akol